# Retrieval-Guided Compositional Image Generation for Cross-Domain Few-Shot Object Detection

### v1

@@ 4.1 Main Comparison Results
CD-FSOD Results. Tab. 1 summarizes the main comparison results on CD-FSOD under 1/5/10 shots across six novel targets. Particularly, we include several competitors: Meta-RCNN [61], TFA w/cos [62], FSCE [63], DeFRCN [64], Distill-cdfsod [65], ViTDeT-FT [66], Detic/Detic-FT [67], DE-ViT [68] as reported in CD-ViTO [1]. In addition, we also report the results of fine-tuned GroundingDINO [28], ETS [27] to compare with our Domain-RAG. Note that both ETS and Domain-RAG build on GroundingDINO but with different augmentation strategies.
@@
RS-FSOD Results. Tab. 2 summarizes the results on the NWPU VHR-10 remote sensing dataset [58] under the 3/5/10/20-shot settings. The dataset is divided into $7$ base classes and $3$ novel classes. The table is split into two parts. In the upper part, we follow the standard RS-FSOD protocol: models are first trained on the base classes and then fine-tuned and evaluated on the novel classes. Under this setting, the base classes contain a sufficient number of annotated samples, and we apply our augmentation strategy on top of the previous state-of-the-art method SEA-FSDet [33], and report the mean Average Precision (mAP) over the 3 novel classes. In the lower part, we explore the dataset in a CD-FSOD setting, where the pretrained model is directly fine-tuned on all 10 classes (both base and novel), each with only a few labeled samples. To ensure comparability with the upper setting, the reported mAP here reflects performance exclusively on the three novel categories.
@@
Meta-RCNN [61] RS-FSOD ResNet-50 20.51 21.77 26.98 28.24 24.38
FsDetView [69] RS-FSOD ResNet-50 24.56 29.55 31.77 32.73 29.65
TFA w/cos [62] RS-FSOD ResNet-50 16.17 20.49 21.22 21.57 19.86
P-CNN [70] RS-FSOD ResNet-50 41.80 49.17 63.29 66.83 55.27
FSOD [71] RS-FSOD ResNet-50 10.95 15.13 16.23 17.11 14.86
FSCE [63] RS-FSOD ResNet-50 41.63 48.80 59.97 79.60 57.50
ICPE [72] RS-FSOD ResNet-50 6.10 9.10 12.00 12.20 9.85
VFA [73] RS-FSOD ResNet-50 13.14 15.08 13.89 20.18 15.57
SAE-FSDet [33] RS-FSOD ResNet-50 57.96 59.40 71.02 85.08 68.36
@@
Camouflaged FSOD Results. Tab. 3 presents the results on the CAMO-FS [34] under 1/2/3/5 shots. All categories in this dataset are treated as novel classes and are further split into a support set and a query set, naturally aligning with the formulation of CD-FSOD. The first two rows in the table report the results of "FS-CDIS-ITL" and "FS-CDIS-IMS", two methods developed from the original CAMO-FS paper. Below that, we include our reproduced baseline using GroundingDINO as the detector, along with the results of our proposed Domain-RAG method built on top of GroundingDINO.
As shown by the results, the large-scale pretrained model, i.e., GroundingDINO, brings a substantial performance boost to this task, improving results from around $7$ to over $65$ mAP. We believe this remarkable advancement will advance the frontier of this field. Moreover, the performance gains introduced by our proposed method over the GroundingDINO baseline remain consistently clear across all shot settings. The consistent success across CD-FSOD, RS-FSOD, and camouflaged FSOD, covering eight challenging and diverse domains, demonstrates that our method serves as a general and effective solution for addressing the gap issue in few-shot object detection.
@@
FS-CDIS-ITL [34] ResNet-101 4.0 7.3 7.5 9.8 7.1
FS-CDIS-IMS [34] ResNet-101 4.5 7.0 7.6 10.4 7.4
@@ 4.2 Comparison with Other Augmentation Methods
To assess the effectiveness of our Domain-RAG framework, we compare it with several strong baselines that are designed for augmenting data for CD-FSOD. Specifically, 1) "Copy-Paste" directly overlays foreground objects onto random COCO backgrounds without considering semantic relevance or compositional integrity. 2) "Foreground Augmentation" attempts to diversify object appearances by inpainting new foregrounds after object removal. This is done by using the category label of each bounding box as a text prompt and applying SDXL-inpaint to generate a new foreground after removing the original object. 3) "Background Augmentation", which we use InstructBLIP [74] to caption the remaining background, and guide SDXL to generate a new background based on this caption after removing the foreground from a target image. To ensure fair comparison, all the augmentation methods use $G=5$. Comparison results are summarized in Tab. 4. The results are reported on CD-FSOD under the 1-shot setting.
Table 4: Comparison of augmentation methods (mAP) on the CD-FSOD benchmark under 1 shot.
@@ 4.3 More Analysis
Figure 3: a) Ablation study on modules, results reported on NEU-DET, 1 shot. (b) Visualization of target image (top row) and generated image (second row). (c) Failure Cases.
@@ Appendix A Technical Appendices and Supplementary Material
Domain-Guided Background Generation. In this phase, the target image and each retrieved image are processed by Redux [14]; their embeddings are combined by a weighted sum (1.0 for the target, 0.8 for the retrieval) without any additional textual prompt. The resulting embedding is fed into the FLUX [9] diffusion model with a guidance scale of 2.5 and 50 sampling steps to synthesize a 1024 × 1024 background image.
Foreground-Background Composition. In the generation stage, we employ the FLUX-Fill [38] model. Given a target image and its background, we supply FLUX-Fill with the source image, an object-mask that excludes the out-painting bounding box, and a prompt embedding extracted from the second-stage background image via the Redux encoder; no additional textual cue is provided and both weights are kept at 1. Because FLUX-Fill struggles with very small inputs, we introduce an adaptive rescaling strategy. For UODD [57], where bounding boxes are tiny, we preserve the aspect ratio and iteratively upsample the image until its longer side exceeds 2048 pixels, then generate a corresponding upsampled mask. Conversely, in ArTaxOr [52], some images exceed 4000 × 3000 pixels and would exhaust GPU memory; whenever either edge is larger than 2800 pixels, we downsample by an integer factor and create a matching mask. The (up- or down-)sampled target image, its mask, and the Redux embedding are then passed to FLUX-Fill. We keep the guidance scale at 30.0 but modulate the overall noise strength to suit each target, with 0.8 for FISH [55] and DIOR [54], 0.9 for ArTaxOr and clipart1k [53], 0.3 for NEU-DET [56] 0.8 for NWPU VHR-10 [58], 0.6 for Camouflage [34] FSOD benchmark, and 0.4 for UODD.
@@
During training, following Grounding DINO [28] and ETS [27], we apply diverse data augmentations, including Mosaic, MixUp, color jittering, random flipping, multi-scale resizing, and cropping, to improve few-shot generalization of both the base GroundingDINO and our Domain-RAG. For evaluation, we follow COCO metrics and disable all augmentations.
@@
Ablation on Text-to-Image Backbones. In this section, we conduct ablation experiments on the background generation module. In the proposed framework, we use the Redux module to fuse features from the retrieval image and the inpainted image, and convert them into a prompt embedding. For the ablation setting, we replace Redux with InstructBLIP and use Diffusion XL to generate backgrounds. Specifically, we use InstructBLIP to extract captions for both the retrieval image and the inpainted image. The resulting captions are then concatenated in the form of "$<caption_{1}>.<caption_{2}>$". Since Diffusion XL uses a CLIP text encoder that cannot handle excessively long texts, we constrain InstructBLIP to generate captions of no more than 20 words, and truncate any prompt exceeding 40 words before feeding it into the model. In this ablation, the background generation strategy is only applied to the domain-guided background generation, while all other components of the pipeline remain unchanged. As shown in Tab. 10, generating richer backgrounds contributes positively to the final generation quality. Our adopted Flux + Redux approach provides more effective supervision, guiding the generation process to produce data that better aligns with the target domain.

### v2

@@ 4.1 Main Comparison Results
CD-FSOD Results. Tab. 1 summarizes the main comparison results on CD-FSOD under 1/5/10 shots across six novel targets. Particularly, we include several competitors: Meta-RCNN [yan2019meta], TFA w/cos [wang2020frustratingly], FSCE [sun2021fsce], DeFRCN [qiao2021defrcn], Distill-cdfsod [xiong2023cd], ViTDeT-FT [li2022exploring], Detic/Detic-FT [zhou2022detecting], DE-ViT [zhang2023detect] as reported in CD-ViTO [fu2024cross]. In addition, we also report the results of fine-tuned GroundingDINO [liu2024grounding], ETS [pan2025enhance] to compare with our Domain-RAG. Note that both ETS and Domain-RAG build on GroundingDINO but with different augmentation strategies.
@@
RS-FSOD Results. Tab. 2 summarizes the results on the NWPU VHR-10 remote sensing dataset [niemeyer2014contextual] under the 3/5/10/20-shot settings. The dataset is divided into $7$ base classes and $3$ novel classes. The table is split into two parts. In the upper part, we follow the standard RS-FSOD protocol: models are first trained on the base classes and then fine-tuned and evaluated on the novel classes. Under this setting, the base classes contain a sufficient number of annotated samples, and we apply our augmentation strategy on top of the previous state-of-the-art method SEA-FSDet [liu2024few], and report the mean Average Precision (mAP) over the 3 novel classes. In the lower part, we explore the dataset in a CD-FSOD setting, where the pretrained model is directly fine-tuned on all 10 classes (both base and novel), each with only a few labeled samples. To ensure comparability with the upper setting, the reported mAP here reflects performance exclusively on the three novel categories.
@@
Meta-RCNN [yan2019meta] RS-FSOD ResNet-50 20.51 21.77 26.98 28.24 24.38
FsDetView [kang2019few] RS-FSOD ResNet-50 24.56 29.55 31.77 32.73 29.65
TFA w/cos [wang2020frustratingly] RS-FSOD ResNet-50 16.17 20.49 21.22 21.57 19.86
P-CNN [cheng2021prototype] RS-FSOD ResNet-50 41.80 49.17 63.29 66.83 55.27
FSOD [fan2020few] RS-FSOD ResNet-50 10.95 15.13 16.23 17.11 14.86
FSCE [sun2021fsce] RS-FSOD ResNet-50 41.63 48.80 59.97 79.60 57.50
ICPE [lu2023breaking] RS-FSOD ResNet-50 6.10 9.10 12.00 12.20 9.85
VFA [han2023few] RS-FSOD ResNet-50 13.14 15.08 13.89 20.18 15.57
SAE-FSDet [liu2024few] RS-FSOD ResNet-50 57.96 59.40 71.02 85.08 68.36
@@
Camouflaged FSOD Results. Tab. 3 presents the results on the CAMO-FS [nguyen2024art] under 1/2/3/5 shots. All categories in this dataset are treated as novel classes and are further split into a support set and a query set, naturally aligning with the formulation of CD-FSOD. The first two rows in the table report the results of "FS-CDIS-ITL" and "FS-CDIS-IMS", two methods developed from the original CAMO-FS paper. Below that, we include our reproduced baseline using GroundingDINO as the detector, along with the results of our proposed Domain-RAG method built on top of GroundingDINO.
As shown by the results, the large-scale pretrained model, i.e., GroundingDINO, brings a substantial performance boost to this task, improving results from around $7$ to over $65$ mAP.
We believe this remarkable advancement will advance the frontier of this field. Moreover, the performance gains introduced by our proposed method over the GroundingDINO baseline remain consistently clear across all shot settings. The consistent success across CD-FSOD, RS-FSOD, and camouflaged FSOD, covering eight challenging and diverse domains, demonstrates that our method serves as a general and effective solution for addressing the gap issue in few-shot object detection.
@@
FS-CDIS-ITL [nguyen2024art] ResNet-101 4.0 7.3 7.5 9.8 7.1
FS-CDIS-IMS [nguyen2024art] ResNet-101 4.5 7.0 7.6 10.4 7.4
@@ 4.2 Comparison with Other Augmentation Methods
To assess the effectiveness of our Domain-RAG framework, we compare it with several strong baselines that are designed for augmenting data for CD-FSOD. Specifically, 1) "Copy-Paste" directly overlays foreground objects onto random COCO backgrounds without considering semantic relevance or compositional integrity. 2) "Foreground Augmentation" attempts to diversify object appearances by inpainting new foregrounds after object removal. This is done by using the category label of each bounding box as a text prompt and applying SDXL-inpaint to generate a new foreground after removing the original object. 3) "Background Augmentation", which we use InstructBLIP [instructblip] to caption the remaining background, and guide SDXL to generate a new background based on this caption after removing the foreground from a target image. To ensure fair comparison, all the augmentation methods use $G=5$. Comparison results are summarized in Tab. 4. The results are reported on CD-FSOD under the 1-shot setting.
Table 4: Comparison of augmentation methods (mAP) on the CD-FSOD benchmark under 1-shot.
@@ 4.3 More Analysis
The Construction of RAG Database In the defined (closed-source) CD-FSOD setting as proposed in CD-ViTO [fu2024cross], COCO serves as the only single-source dataset for training, while other datasets (ArTaxOr, Clipart1k, DIOR, DeepFish, NEU-DET, UODD) are treated as unseen targets. Using COCO as the RAG database brings two key advantages: (1) It does not introduce any extra data beyond the default setting, ensuring the fairness of comparison; (2) COCO provides diverse and general-domain backgrounds that better cover novel domain scenarios.
Furthermore, we conducted additional experiments using different database options, including COCO with reduced category numbers, NEU-DET (non-general-domain), and miniImageNet.
Table 5: Effect of different database choices on Domain-RAG performance.
DataBase ArTaxOr Clipart1k DIOR FISH NEU-DET UODD Avg
Base (GroundingDINO) 26.3 55.3 14.8 36.4 9.3 15.9 26.3
COCO-1class 50.1 55.0 15.7 36.6 12.0 16.0 30.9
COCO-5classes 51.0 55.1 16.6 36.7 11.9 17.5 31.5
COCO-20classes 53.0 56.2 16.2 37.0 12.2 18.9 32.3
COCO-80classes (Ours) 57.2 56.1 18.0 38.0 12.1 20.2 33.6
NEU-DET 49.8 55.2 16.4 37.0 12.0 16.1 31.1
miniImageNet 55.6 53.2 15.6 38.0 14.0 16.2 32.1
From the results summarized in Table 5, we observe that: (1) broader category coverage consistently improves performance; (2) general-domain databases such as COCO outperform specific-domain ones like NEU-DET; and (3) although miniImageNet can serve as an alternative database, it performs slightly worse than COCO due to its larger foreground regions and less diverse backgrounds. These findings demonstrate that our Domain-RAG consistently enhances the base GroundingDINO across all benchmarks, validating the robustness and effectiveness of our approach.
@@
Figure 3: (a) Ablation study on modules, results reported on NEU-DET, 1 shot. (b) Visualization of target image (top row) and generated image (second row). (c) Failure Cases.
@@ 5 Conclusion
6 Acknowledgment
This work was supported by the Science and Technology Commission of Shanghai Municipality (No. 24511103100).The authors gratefully thank the organization for their support and resources.
@@ Appendix A Technical Appendices and Supplementary Material
Domain-Guided Background Generation. In this phase, the target image and each retrieved image are processed by Redux [fluxredux2024]; their embeddings are combined by a weighted sum (1.0 for the target, 0.8 for the retrieval) without any additional textual prompt. The resulting embedding is fed into the FLUX [FLUX] diffusion model with a guidance scale of 2.5 and 50 sampling steps to synthesize a 1024 × 1024 background image.
Foreground-Background Composition. In the generation stage, we employ the FLUX-Fill [fluxfill2024] model. Given a target image and its background, we supply FLUX-Fill with the source image, an object-mask that excludes the out-painting bounding box, and a prompt embedding extracted from the second-stage background image via the Redux encoder; no additional textual cue is provided and both weights are kept at 1. Because FLUX-Fill struggles with very small inputs, we introduce an adaptive rescaling strategy. For UODD [jiang2021underwater], where bounding boxes are tiny, we preserve the aspect ratio and iteratively upsample the image until its longer side exceeds 2048 pixels, then generate a corresponding upsampled mask. Conversely, in ArTaxOr [GeirArTaxOr], some images exceed 4000 × 3000 pixels and would exhaust GPU memory; whenever either edge is larger than 2800 pixels, we downsample by an integer factor and create a matching mask. The (up- or down-)sampled target image, its mask, and the Redux embedding are then passed to FLUX-Fill. We keep the guidance scale at 30.0 but modulate the overall noise strength to suit each target, with 0.8 for FISH [saleh2020realistic] and DIOR [li2020object], 0.9 for ArTaxOr and clipart1k [inoue2018cross], 0.3 for NEU-DET [song2013noise] 0.8 for NWPU VHR-10 [niemeyer2014contextual], 0.6 for Camouflage [nguyen2024art] FSOD benchmark, and 0.4 for UODD.
@@
During training, following Grounding DINO [liu2024grounding] and ETS [pan2025enhance], we apply diverse data augmentations, including Mosaic, MixUp, color jittering, random flipping, multi-scale resizing, and cropping, to improve few-shot generalization of both the base GroundingDINO and our Domain-RAG. For evaluation, we follow COCO metrics and disable all augmentations.
@@
Ablation on Text-to-Image Backbones. In this section, we conduct ablation experiments on the background generation module. In the proposed framework, we use the Redux module to fuse features from the retrieval image and the inpainted image, and convert them into a prompt embedding. For the ablation setting, we replace Redux with InstructBLIP and use Diffusion XL to generate backgrounds. Specifically, we use InstructBLIP to extract captions for both the retrieval image and the inpainted image. The resulting captions are then concatenated in the form of $caption_{1}>.<caption_{2}>$". Since Diffusion XL uses a CLIP text encoder that cannot handle excessively long texts, we constrain InstructBLIP to generate captions of no more than 20 words, and truncate any prompt exceeding 40 words before feeding it into the model. In this ablation, the background generation strategy is only applied to the domain-guided background generation, while all other components of the pipeline remain unchanged. As shown in Tab. 11, generating richer backgrounds contributes positively to the final generation quality. Our adopted Flux + Redux approach provides more effective supervision, guiding the generation process to produce data that better aligns with the target domain.